### v3

@@ 5 Results
Table 2: Results of CSEGG for various continual learning methods applied on the two SGG backbones (SGTR and TCNN) in Learning Scenarios 1 and 2. See Sec. 3.2 for continual learning baselines. See Sec. 3.3 for evaluation metrics. The higher the evaluation metrics, the better. The best are in bold.
@@ 5.1 RAS outperforms all the CSEGG baselines in Scenarios 1 and 2
Table 3: Results at Task 5 in Learning Scenario 1 when sampling techniques are applied to long-tailed distribution data. See Sec. 3.2 for the introduction to techniques used for long-tailed distributions. We copy the results of our RAS from Tab. 2 for easy comparisons. The best results are in bold.
Figure 5: Results in Learning Scenario 3. See Sec. 3.3 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning methods. Line types denote the IoU thresholds.
The results for Avg. R, F, mR, mF, FWT, and BWT in learning scenarios 1 (S1) and 2 (S2) are presented in Tab. 2. Our observations align with established research in continual learning, especially in image classification and object detection: regardless of the SGG architectures, over both learning scenarios, Naive consistently performs the worst, showcasing significant catastrophic forgetting. Rehearsal-based methods, such as Replay@10% and Replay@20%, outperform techniques like EWC and PackNet. However, none of them surpass our RAS.
Moreover, compared to Replay@20%, RAS achieves superior performance while requiring a smaller constant storage size, equivalent to storing 15% exemplary images from the training set in the Replay methods. This storage size is determined by the image dimensions of 800 by 600 pixels and the number of model parameters in RAS. It is also worth noting that RAS only requires storing the old model snapshot, the triplet labels in $U_{t}$, and the image generative model, eliminating privacy concerns associated with storing original images.
In Learning Scenario 2 (S2), the task involves classifying both new objects and new relationships, significantly escalating the level of difficulty compared to S1. As evident from Tab. 2, all CSEGG models, including Replay@100%, exhibit a decline in overall performance when compared to the upper bound Joint. Although our RAS achieves the leading performance among all the baselines, its performance is still far from Joint. This underscores the persistent challenge posed by S2 in the context of CSEGG. Future work should explore new approaches to address the complexities inherent in this scenario.
@@
As the knowledge carried forward is important for the subsequent tasks, we also permuted the task sequences and explored their role in CSEGG performances. Aligning with the existing literature [54], we found a prominent effect of task sequences in CSEGG (Fig. S3 and Sec. 0.A.4.1). To gain a qualitative understanding of CSEGG performances, we provide visualization results of the predicted scene graphs on example images over tasks for all the CSEGG baselines in Scenario 1 (Fig. S5 and Sec. 0.A.5.1) and Scenario 2 (Fig. S6 and Sec. 0.A.5.2).
@@ 5.2 CSEGG Models Can Generalize in Unknown Scenes
Fig. 5 illustrates the generalization results for detecting unknown objects and classifying known relationships among these objects in Learning Scenario 3 (S3). In Fig. 5 (a), an increasing trend in Gen Rbbox is observed with the increasing task number for all CSEGG methods, indicating improved generalization in detecting unknown objects. Notably, even with minimal training in Task 1, all CSEGG methods propose 23% reasonable object regions with IoU = 0.7, showcasing the SGTR model’s ability to generalize “objectness". As expected, with an increase in IoU threshold from 0.3 to 0.7, we found that Gen Rbbox decreases due to fewer bounding boxes being considered correct. Moreover, we also compared the generalization performance in object detection between Replay@10% and Naive. Contrary to previous observations in S1 and S2 (Tab. 2), we found that Replay@10% show a decline in Gen Rbbox, possibly due to a fixed number of detected object bounding boxes output by CSEGG methods. Similarly, our RAS exhibits a reduced G Rbbox compared to Replay@10% and Naive, likely for the same underlying reason.
In Fig. 5 (b), Replay@10% outperforms Naive in Gen R when considering correctly detected unknown object locations, emphasizing that minimizing forgetting in continual learning enhances the SGTR model’s overall relationship generalization in unknown scene understanding. However, the performance of Replay@10% is still inferior to our RAS method; implying that our RAS is more proficient in generalizing to classify relationships among unknown objects. Interestingly, we also noted that even with minimal training in Task 1, all the CSEGG methods achieve 45% recall of known relationships among unknown objects, demonstrating the SGTR model’s ability to generalize “relationship". Visualization examples, when CSEGG models can generalize to recognize relationships, are presented in Fig.S7 and Sec.0.A.5.3.
@@ 5.3 Ablation Studies on Our RAS Reveal Key Design Insights
Table 4: Ablation results of our RAS on learning scenario S1 reveals key design insights. This table presents the results of ablation studies conducted to identify key components of our method, as discussed in Sec. 5.3. Results in Avg.R, F, mR, mF are reported after the last task in the scenario S1.
We introduce our default method designs in Sec. 4. Here, we vary the components in our RAS to reveal key design insights.
We propose a context checker in RAS. Here, we conduct an ablation by removing this module. Triplet labels are randomly selected and combined for text prompts. In A1 of Tab. 4, we observe a performance decrease of approximately 2% across all evaluation metrics, compared with our RAS. This suggests that generating images adhering to real-world context rules is crucial for replays. The lower performance may be attributed to the challenge of generating good-quality out-of-context images for Stable Diffusion Models and the potential domain differences affecting the SGG model $M_{t}$ in predicting out-of-context SGG notations.
@@
Lastly, we investigate the impact of generating $\gamma$ images using the same text prompt in RAS. We vary $\gamma$ from 2 to 8 and present their corresponding model performances in A4-6 of Tab. 4. As expected, the SGG model’s performance improves with higher $\gamma$, indicating that generating more images increases sample diversity for rehearsals and hence, it enhances CSEGG performance. We set $\gamma=10$ as the default; however, in practical scenarios with ample compute resources, dynamically synthesizing more images could further improve CSEGG performances. This also highlights the advantage of our RAS, as it can dynamically generate numerous images for replays without expanding its storage usage.
@@ 6 Discussion
In the dynamic world, adapting scene graph generation (SGG) models to new objects and relationships poses challenges. Despite progress in SGG and continual learning, there is still a gap in understanding Continual Scene Graph Generation (CSEGG). We address this by operationalizing CSEGG, and introducing benchmarks, datasets, and evaluation protocols. Our study explores three learning scenarios, analyzing continual object detection and relationship classification in long-tailed class-incremental settings for CSEGG baselines. Our findings show that integrating sampling methods with CSEGG baselines to address long-tailed distributions moderately eliminates forgetting; however, a large performance gap between current CSEGG baselines and the joint training upper bound persists. To address CSEGG challenges, we propose RAS, a Replays via Analysis by Synthesis method. RAS parses previous task scene graphs into triplet labels for diverse in-context scene graph reconstruction. Based on these re-compositional context-congruent scene graphs, RAS synthesizes images with Stable Diffusion models for replays. Unlike other image replay methods, RAS stores only triplet labels and the model snapshot, maintaining constant memory usage and preserving privacy. Extensive experiments demonstrate our RAS’s superior performance over current CSEGG baselines in knowledge transfers and reducing forgetting. Interestingly, our RAS model is also capable of generalizing to classify known relationships among unfamiliar objects. We discuss ethical concerns in the Sec.0.A.6.
7 Acknowledgement
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), its NRFF award NRF-NRFF15-2023-0001, Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR. The authors declare that they have no competing interests. The authors would like to thank Stan Weixian Lei, Difei Gao, and Mike Shou for their feedback and suggestions.
Appendix 0.A Appendix
0.A.1 Introduction to Three Learning Scenarios
@@
0.A.1.1 Scenario 1 (S1): Relationship Incremental Learning
While existing continual object detection literature focuses on incrementally learning object attributes [37, 61, 8, 59, 51, 39], incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene. To uncover contextual information and go beyond studies of object attributes, we introduce this scenario where new relationship predicates $p_{k}$ are incrementally added in each task (Fig. 2 S1). There are 5 tasks in S1. To simulate the naturalistic settings where the frequency of relationship distribution is often long-tailed, we randomly and uniformly sample relationship classes from head, body and tail categories in Visual Genome [23], and form a set of 10 relationship classes for each task. Thus, the relationships within a task are long-tailed; and the number of relationships from the head categories of each task is of the same scale. To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task (see Sec. 0.A.1.4 for the design motivation). The same reasoning applies in S2 and S3. Example relationship classes from each task and their distributions are provided in Fig. S1.
@@ 0.A.1.1 Scenario 1 (S1): Relationship Incremental Learning
0.A.1.2 Scenario 2 (S2): Scene Incremental Learning
@@
0.A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
@@
0.A.1.4 Design Motivations for Overlapping Images and Non-overlapping Labels across Tasks
The design of such images and label splits over tasks aligns with human learning scenarios where a parent teaches the baby to recognize different toys and objects in the bedroom. Though the baby is exposed to the same bedroom scenes multiple times, the parent only teaches the baby to detect and recognize one object at a time in a continual learning setting. In the future, we will expand our studies to cases where the SGG models learn from non-overlapping sets of training images for each task.
0.A.2 Implementation and Training Details
@@
Figure S2: Introduction to backbone SGG models and continual learning baselines. We use Scene graph Generation TRansformer (SGTR) [28] as the backbone SGG model (Sec. 3.2). SGTR consists of four modules indicated by each blue box. Arrows indicate the signal flows among modules. (b) Four continual learning baselines are listed: EWC [22], Replay [48], Naive (Sec. 3.2) and PackNet [38](Sec. 3.2). $\theta^{*}_{A}$ denotes the optimal network parameters after learning on task A. The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
0.A.2.1 SGTR Backbone
@@
For SGTR as depicted in (Fig. S2 (a)), the approach uniquely formulates the task as a bipartite graph construction problem. Starting with a scene image ($I_{i}$), SGTR utilizes a 2D-CNN and transformer-based encoder to extract image features. These features are then incorporated into a transformer-based decoder, predicting object and subject nodes ($O_{i}$). Predicate nodes ($R_{i}$) are formed based on both image features and object node features, and a bipartite graph ($G_{i}$) is constructed to represent the scene collectively. The correspondence between object nodes ($o_{i}$) and predicate nodes ($r_{k}$) is established using the Hungarian matching algorithm [24]. Experimental results are based on the average over three runs, and the implementation leverages public source codes from [28] and [59] with default hyperparameters.
@@
The SGTR is trained in two stages in a supervised manner. In stage 1, only object detection losses in DETR is [6] applied on $O_{i}$. In stage 2, only predicate entity prediction loss is applied on $R_{i}$, which can be further decomposed into L1 and GIOU losses for object/subject/predicate localization [46] and cross-entropy loss for object/subject/predicate classification. In learning scenario S1, we skip Stage 1, and directly load pre-trained weights of DETR for object detection on the entire training set of Visual Genome [28]. In stage 2 of S1, we freeze the feature extractor, and fine-tune the rest parts of SGTR for predicate entity predictions. As only relationship classes are incrementally introduced in S1, we freeze the entire weights of DETR for detecting all the objects in the scene over tasks. However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions [28]. Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Sec. S4 and Sec. 0.A.4.2). Thus, the forgetting observed in S1 could only be attributed to incremental relationship learning. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the feature extractor, pre-trained on ImageNet [12], and fine-tune the entire DETR [6] over all the tasks. Stage 2 of S2 and S3 is the same as S1.

### v4

@@ 5 Results
Table 2: Results of CSEGG for various continual learning methods applied on the two SGG backbones (SGTR and TCNN) in Learning Scenarios 1 and 2. See Sec. 3.2 for continual learning baselines. See Sec. 3.3 for evaluation metrics. The higher the evaluation metrics, the better. The best are in bold. * means the experiment is still running, we will report the results in the final version.
@@ 5.1 RAS outperforms all the CSEGG baselines in Scenarios 1 and 2
Table 3: Results at Task 5 in Learning Scenario 1 when sampling techniques are applied to long-tailed distribution data. See Sec. 3.2 for the sampling techniques on long-tailed distributions. We copy the results of our RAS from Tab. 2 for easy comparisons. The best results are in bold.
Figure 4: Results in Scenario 3. See Sec. 3.3 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning methods. Line types denote the IoU thresholds.
The results for Avg. R, F, mR, mF, FWT, and BWT in learning scenarios 1 (S1) and 2 (S2) are presented in Tab. 2. Our observations align with established research in continual learning, especially in image classification and object detection: regardless of the SGG architectures, over both learning scenarios, Naive consistently performs the worst, showcasing significant catastrophic forgetting. Replay-based methods, such as Replay@10% and Replay@20%, outperform techniques like EWC and PackNet. However, none of them surpass our RAS.
RAS achieves superior performance compared to Replay@20% ($\sim$2 Gb) while requiring less storage ($\sim$1.2 Gb), equivalent to storing 15% exemplary images. This storage efficiency is due to only needing to store the old model snapshot, triplet labels in $U_{t}$, and the image generation model, thus avoiding privacy concerns. Also, RAS outperforms RAS_GT (Tab. 2), indicating that decomposing scene graphs into smaller, more diverse ones, along with more comprehensible prompts, is more effective than storing the ground truth scene graphs and directly using them for image generation.
In Learning Scenario 2 (S2), the task involves classifying both new objects and new relationships, significantly escalating the level of difficulty compared to S1. As evident from Tab. 2, all CSEGG models, including Replay@100%, exhibit a decline in overall performance when compared to the upper bound Joint. Although our RAS achieves the leading performance among all the baselines, its performance is still far from Joint. This underscores the persistent challenge posed by S2 in the context of CSEGG. Future work should explore new approaches to address this gap.
@@
We explored the impact of task sequence permutations on CSEGG performance, finding an effect consistent with existing literature [63] (Fig. S6; Sec. A.6.1). We also observed that fine-tuning DETR in S1 has minimal impact on forgetting, indicating that any forgetfulness in S1 is solely due to relationship incremental learning (Fig. S7; Sec. A.6.2). Moreover, to gain qualitative insights, we provide visualizations of predicted scene graphs for all CSEGG baselines in Scenario 1 (Fig. S8 and Sec. A.7.1) and Scenario 2 (Fig. S9 and Sec. A.7.2).
@@ 5.2 CSEGG Models Can Generalize in Unknown Scenes
Fig. 4 illustrates the generalization results for detecting unknown objects and classifying known relationships among these objects in Learning Scenario 3 (S3). In Fig. 4 (a), an increasing trend in Gen Rbbox is observed with the increasing task number for all CSEGG methods, indicating improved generalization in detecting unknown objects. Notably, even with minimal training in Task 1, all CSEGG methods propose 23% reasonable object regions with threshold IoU = 0.7, showcasing the SGTR model’s ability to generalize to locate “objectness". As expected, with an increase in IoU threshold from 0.3 to 0.7, we found that Gen Rbbox decreases due to fewer bounding boxes being considered correct. Moreover, we also compared the generalization performance in object detection between Replay@10% and Naive. Contrary to previous observations in S1 and S2 (Tab. 2), we found that Replay@10% show a decline in Gen Rbbox, possibly due to a fixed number of detected object bounding boxes output by CSEGG methods. Similarly, our RAS exhibits a reduced G Rbbox compared to Replay@10% and Naive, likely for the same underlying reason.
In Fig. 4 (b), Replay@10% outperforms Naive in Gen R when considering correctly detected unknown object locations, emphasizing that minimizing forgetting in continual learning enhances the SGTR model’s overall relationship generalization in unknown scene understanding. However, the performance of Replay@10% is still inferior to our RAS method; implying that our RAS is more proficient in generalizing to classify relationships among unknown objects. Interestingly, we also noted that even with minimal training in Task 1, all the CSEGG methods achieve 45% recall of known relationships among unknown objects, demonstrating the SGTR model’s ability to generalize to classify “relationship". Visualization examples, when CSEGG models can generalize to recognize relationships, are presented in Fig. S10 and Sec. A.7.3.
@@ 5.3 Ablation Studies on Our RAS Reveal Key Design Insights
Table 4: Ablation results of our RAS on learning scenario S1 reveals key design insights. This table presents the results of ablation studies conducted to identify key components of our method, as discussed in Sec. 5.3. Results in Avg.R, F, mR, mF are reported after the last task in Scenario S1.
We introduce our default method designs in Sec. 4. Here, we vary the components in our RAS to reveal key design insights. We propose a context checker in RAS. Here, we conduct an ablation by removing this module. Triplet labels are randomly selected and combined for text prompts. In A1 of Tab. 4, we observe a performance decrease of approximately 2% across all evaluation metrics, compared with our RAS. This suggests that generating images adhering to real-world context rules is crucial for replays. The lower performance may be attributed to the challenge of generating good-quality out-of-context images for Stable Diffusion Models and the potential domain differences affecting the SGG model $M_{t}$ in predicting out-of-context SGG notations.
@@
Lastly, we investigate the impact of generating $\gamma$ images using the same text prompt in RAS, varying $\gamma$ from 2 to 8 (Tab. 4, A4-6). As expected, performance improves with higher $\gamma$, showing that increased sample diversity enhances CSEGG performance. With ample computing resources, dynamically synthesizing more images could further improve performance. This highlights RAS’s advantage in generating numerous images for replays without expanding storage usage.
@@ 6 Discussion
In the dynamic world, adapting scene graph generation (SGG) models to new objects and relationships poses challenges. Despite progress in SGG and continual learning, there is still a gap in understanding Continual Scene Graph Generation (CSEGG). We address this by operationalizing CSEGG, and introducing benchmarks, datasets, and evaluation protocols. Our study explores three learning scenarios, analyzing continual object detection and relationship classification in long-tailed class-incremental settings for CSEGG baselines. Our findings show that integrating sampling methods with CSEGG baselines to address long-tailed distributions moderately eliminates forgetting; however, a large performance gap between current CSEGG baselines and the joint training upper bound persists. To address CSEGG challenges, we propose RAS, a Replays via Analysis by Synthesis method. RAS parses previous task scene graphs into triplet labels for diverse in-context scene graph reconstruction. Based on these re-compositional context-congruent scene graphs, RAS synthesizes images with Stable Diffusion models for replays. Unlike other image replay methods, RAS stores only triplet labels and the model snapshot, maintaining constant memory usage and preserving privacy. Extensive experiments demonstrate our RAS’s superior performance over current CSEGG baselines in knowledge transfers and reducing forgetting. Interestingly, our RAS model is also capable of generalizing to classify known relationships among unseen objects.
Moving forward, there are several key avenues for future research. First, our current endeavors focus on tackling CSEGG problems from static images in an Independent and Identically Distributed (i.d.d) manner, diverging from how humans learn from video streams. Future research can look into CSEGG problems on video SGG datasets. Second, our plans also involve expanding the set of continual learning baselines and integrating more long-tailed distribution sampling techniques. Third, we aim to construct a synthetic SGG dataset to systematically quantify the aspects of SGG that influence continual learning performance under controlled conditions. In RAS, SGG annotations for synthesized images in the replay buffer are predicted by the preceding SGG model, which can lead to error propagation across training iterations. In the future work, integrating a generative model with fine-grained control signals (such as bounding boxes and captions) [36] could provide more precise supervision, potentially mitigating these accumulated errors and further enhancing the performance of our approach. Although the CSEGG method holds promise for many downstream applications like monitoring systems, medical imaging, and autonomous navigation, we should also be aware of its misuse in privacy, data biases, fairness, security concerns, and misinterpretation (see Sec. A.8 for an expanded discussion). We invite the research community to join us in expanding and updating the safe use of CSEGG benchmarks, thereby fostering its advancements in research and technology.
Acknowledgement
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), and its NRFF award NRF-NRFF15-2023-0001. We also acknowledge Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), Startup Grant from Nanyang Technological University, and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR.
Appendix A Appendix
A.1 Introduction to Three Learning Scenarios
@@
A.1.1 Scenario 1 (S1): Relationship Incremental Learning
While existing continual object detection literature focuses on incrementally learning object attributes [43, 70, 8, 68, 60, 45], incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene. To uncover contextual information and go beyond studies of object attributes, we introduce this scenario where new relationship predicates $p_{k}$ are incrementally added in each task (Fig. 2 S1). There are 5 tasks in S1. To simulate the naturalistic settings where the frequency of relationship distribution is often long-tailed, we randomly and uniformly sample relationship classes from head, body and tail categories in Visual Genome [25], and form a set of 10 relationship classes for each task. Thus, the relationships within a task are long-tailed; and the number of relationships from the head categories of each task is of the same scale. To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task
The design of such images and label splits over tasks aligns with human learning scenarios where a parent teaches the baby to recognize different toys and objects in the bedroom. Though the baby is exposed to the same bedroom scenes multiple times, the parent only teaches the baby to detect and recognize one object at a time in a continual learning setting. In the future, we will expand our studies to cases where the SGG models learn from non-overlapping sets of training images for each task. The same reasoning applies in S2 and S3. Example relationship classes from each task and their distributions are provided in Fig. S1.
@@ A.1.1 Scenario 1 (S1): Relationship Incremental Learning
A.1.2 Scenario 2 (S2): Scene Incremental Learning
@@
A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
@@
Figure S2: Data Statistics for all Learning Scenarios: (a) Distribution of objects in the entire training set of Visual Genome during Stage 1 of S1. (b) Distribution of relationships during Stage 2 for each task in S1. (c) Distribution of objects during Stage 1 for each task in S2. (d) Distribution of relationships during Stage 2 for each task in S2. (e) Distribution of objects during Stage 1 for each task in S3. (f) Distribution of relationships during Stage 2 for each task in S3. The numbers in brackets in the legend in (b-f) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
A.2 Data Statistics
In this section, we provide various types of data statistics for all three learning scenarios. Specifically, we present statistics regarding the number of images, objects, and relationships involved in each task of each learning scenario. This information is provided in Fig. S2.
A.3 Implementation and Training Details
@@
Figure S3: Introduction to backbone SGG models and continual learning baselines. We use Scene graph Generation TRansformer (SGTR) [32] as the SGG backbone (Sec. 3.2). SGTR consists of four modules indicated by each blue box. Arrows indicate the signal flows among modules. (b) Four continual learning baselines are listed: EWC [24], Replay [57], Naive (Sec. 3.2) and PackNet [44](Sec. 3.2). $\theta^{*}_{A}$ denotes the optimal network parameters after learning on task A. The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
Figure S4: Detailed Schematic of three CSEGG Baselines. In this schematic, we show the operations of each baseline at tasks $t-1$, $t$, $t+1$. (a) shows the schematic for exemplar-based methods like replay and RAS. (b) shows the schematic for EWC. (c) shows the schematic for PackNet. Here, $D_{i}$ refers to the dataset of task $i$. $E_{i}$ denotes the exemplar set for replays at task $i$. $W_{i}$ is the weights of $M_{i}$ at task $i$.
Figure S5: Visualizations of example generated images given the text prompt from our RAS. Each row presents two examples of generated images. Given the input text prompt on the left, the generated image is displayed on the right. These visualizations provide qualitative validations into the capability of RAS to produce diverse and contextually relevant images based on our designed text prompts.
A.3.1 SGTR Backbone
@@
For SGTR in Fig. S3 (a), the approach uniquely formulates the task as a bipartite graph construction problem. Starting with a scene image ($I_{i}$), SGTR utilizes a 2D-CNN and transformer-based encoder to extract image features. These features are then incorporated into a transformer-based decoder, predicting object and subject nodes ($O_{i}$). Predicate nodes ($R_{i}$) are formed based on both image features and object node features, and a bipartite graph ($G_{i}$) is constructed to represent the scene collectively. The correspondence between object nodes ($o_{i}$) and predicate nodes ($r_{k}$) is established using the Hungarian matching algorithm [26]. Experimental results are based on the average over three runs, and the implementation leverages public source codes from [32] and [68] with default hyperparameters.
@@
The SGTR is trained in two stages in a supervised manner. In stage 1, only object detection losses in DETR is [6] applied on $O_{i}$. In stage 2, only predicate entity prediction loss is applied on $R_{i}$, which can be further decomposed into L1 and GIOU losses for object/subject/predicate localization [54] and cross-entropy loss for object/subject/predicate classification. In learning scenario S1, we skip Stage 1, and directly load pre-trained weights of DETR for object detection on the entire training set of Visual Genome [32]. In stage 2 of S1, we freeze the feature extractor, and fine-tune the rest parts of SGTR for predicate entity predictions. As only relationship classes are incrementally introduced in S1, we freeze the entire weights of DETR for detecting all the objects in the scene over tasks. However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions [32]. Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Fig. S7 and Sec. A.6.2). Thus, the forgetting observed in S1 could only be attributed to incremental relationship learning. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the feature extractor, pre-trained on ImageNet [13], and fine-tune the entire DETR [6] over all the tasks. Stage 2 of S2 and S3 is the same as S1.